OpenStudy (anonymous):

Raise 10666369614354225120 to 16479139101483902597. Take the result and divide by 18074899840695706733. Take the rest of that division. Encode it to hex F__-_C__-D__3-58__

OpenStudy (anonymous):

lol that big number

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

i know

OpenStudy (anonymous):

10666369614354225120 = 10666369614354220000 + 320x16 = 16x(20x16 + 1066636961435422x625) = 16 x (16 x 20 + 16 x 39 x 1066636961435422 + 1066636961435422) = 16 x (16 x 20 + 16 x 39 x 1066636961435422 + 16 x 338 x 106663696143 + 14 * 106663696143) = .... 10666369614354225120 two last hex digit is E0 16479139101483900000 + 2597 etc...